A Company manufactures a product A. The company estimates the cost function for the total costs.

A Company manufactures a product A.  The company estimates the cost function for the total costs. The cost driver is number of units.  The following  informations were collected:

 

Month                     Units                                               Total Costs

January                     3,560                                             $242,400

February                   3,800                                             $252,000

March                       4,000                                             $260,000

April                         3,600                                             $244,000

May                          3,200                                             $228,000

June                          3,040                                             $221,600

 

Compute a cost function using the high-low method.
Answer #1

TCHA =Total cost of high activity, HAU=Highest activity unit

·         Variable Cost = (TCHATotal Cost of Low Activity) / (HAU - Lowest Activity Unit) = (260000-220000) / (4000-3040) = 40.625

·         Total Cost = (VC * Units Produced) + Total Fixed Cost

·         Fixed cost = Highest activity cost – (Variable cost per unit x Highest activity units) = 260000-(40.625*4000) = 97500


answered by: nonolll22